[Jamie / Suresh / Kiwamu]
The in-vac work is ongoing.
Before we run out our energy we are posting this entry to briefly report the current status.
- (done) BS earthquake stop adjustment.
- (done) PRM earthquake stop adjustment
- (done) MC spot position check => They are almost the same within 10 %.
- (done) Injection and alignment of the ABSL laser to make the beam brighter in the vertex region.
- (done) POY => We repositioned an in-vac steering mirror to get the POY beam hitting the center of the steering mirror.
It's now coming out from the chamber.
- (done) IPANG => realigned two mirrors on the ETMY table to get the IPANG out from the chamber. Now it's reaching the ETMY optical table.
It needs a final touch before we pump down. We revisited it later in the night after realigning the IFO and it is well aligned now.
- (done) POP => We have aligned the ABSL laser injected from the AS port to reach the REFL camera. We turned it up to max power of 300mW and used it as a substitute for the PRC beam.
Even this was not enough to see anything in the POP beam path after the PR2 (tip-tilt). So we used a green beam from the Y-arm as a guide of the POP beam path because the ABSL (POP) beam was too dim to work with.
We placed a lens and a CCD camera to detect the green and then blocked the Y-green. It was then possible to see the ABSL-POP beam in the CCD camera. The lens and the CCD are markers for this beam.
Do not remove these markers unless absolutely necessary.
-(done) POX => We located the ABSL (mimicking POX) beam on the POXM1 mirror and adjusted the mirror to ensure that the beam exits at the right height and a convenient location on the POX table.
- (0%) OSEM mid-range adjustment
- (0%) IPPOS
- (0%) oplev re-alignment